# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

### v1

@@ Abstract
Existing video captioning benchmarks and models lack coherent representations of causal-temporal narrative, which is sequences of events linked through cause and effect, unfolding over time and driven by characters or agents. This lack of narrative restricts models’ ability to generate text descriptions that capture the causal and temporal dynamics inherent in video content. To address this gap, we propose NarrativeBridge, an approach comprising of: (1) a novel Causal-Temporal Narrative (CTN) captions benchmark generated using a large language model and few-shot prompting, explicitly encoding cause-effect temporal relationships in video descriptions, evaluated automatically to ensure caption quality and relevance; and (2) a dedicated Cause-Effect Network (CEN) architecture with separate encoders for capturing cause and effect dynamics independently, enabling effective learning and generation of captions with causal-temporal narrative. Extensive experiments demonstrate that CEN is more accurate in articulating the causal and temporal aspects of video content than the second best model (GIT): 17.88 and 17.44 CIDEr on the MSVD and MSR-VTT datasets, respectively. The proposed framework understands and generates nuanced text descriptions with intricate causal-temporal narrative structures present in videos, addressing a critical limitation in video captioning. For project details, visit https://narrativebridge.github.io/.
@@ 1 Introduction
Video captioning aims to generate textual descriptions that capture the visual information and temporal dynamics in videos [48]. Research has primarily focused on developing new methods to improve the accuracy of video captioning models on well-established benchmark datasets, MSR-VTT [55] and MSVD [7]. State-of-the-art (SOTA) approaches [19, 44, 56, 32, 51, 9] proposed new architectures to better align the generated captions with the provided ground-truth. However, these efforts enhance models’ accuracy on existing evaluation criteria, without modifying the benchmark datasets or their ground truth annotations to address the lack of coherent representations of causal-temporal narrative - construction and interpretation of a sequence of events linked through cause and effect, unfolding over time and space, and often driven by entities (characters or agents) acting with intention [52].
As discussed by Wilkens et al. [52], causal relationships [16] are scenarios where one event (cause) directly influences the occurrence of another event (effect), while temporal understanding [4] involves recognizing the chronological order of events, a crucial component in establishing causality. Existing ground-truth captions in popular benchmarks lack causal, temporal and narrative information. Figure 1 illustrates the importance of causal-temporal narrative on the MSR-VTT dataset [55]. In the input video sequence, a car is first shown driving recklessly through an open field and flipping over, which represents the cause event. This is then followed by a scene showing the severely damaged car, and subsequently, a group of guys starting to play beer pong, which represents the effect event. The causal-temporal narrative would connect these events in a temporal order, linking the reckless driving and resulting car crash to the group’s subsequent behavior of playing beer pong. This example aligns with the definition of narrative provided by Wilkens et al. [52], which emphasizes the role of cause-effect relationships, the perception of narrativity in videos, and highlights the temporal sequence of events that is crucial for understanding the causal-temporal narrative. Figure 1 highlights the limitations of the original captions in existing benchmark datasets, such as MSR-VTT [55]. The original captions focus on isolated events or actions, such as "a car flipping over" or "a monster truck flips on its side," lacking contextual narrative and causal-temporal relationships between events. Consequently, models trained on these captions suffer from the same limitations.
@@
To bridge this gap, we introduce NarrativeBridge, a novel framework encompassing a new benchmark and architecture tailored for causal-temporal narrative learning in video captioning. Our Causal-Temporal Narrative (CTN) captions benchmark leverages a large language model (LLM) and few-shot prompting to generate enhanced video descriptions that explicitly encode cause-effect relationships and temporal sequences, as exemplified in Figure 1. Our CTN captions benchmark, generated using the approach detailed in Section 3, establishes a clear connection between the cause (reckless driving) and the effect (damaged car and subsequent behavior of the group). By generating a single, coherent caption that encapsulates the causal-temporal narrative of the video, our CTN captions benchmark enables models to better understand and articulate the causality, sequence, and significance of events within the broader video context [52]. This approach addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions. To ensure the quality and relevance of the generated captions, we employ an automatic evaluation framework that compares the CTN captions with the video content, keeping or discarding them based on a score threshold. This approach addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions.
@@
Existing SOTA video captioning methods that use different architectures such as LSTM [32], GNN [17], and Transformer [51], struggle to effectively learn causal-temporal narrative from the CTN captions. These architectures are designed to capture the overall semantics in videos but lack dedicated mechanisms to explicitly model the cause-effect relationships and temporal sequences. As a result, the captions generated by these methods fail to articulate the complex causal-temporal dynamics in the videos, as demonstrated in results (Section 4.4). To address this challenge, we propose the Cause-Effect Network (CEN), an architecture specifically designed to capture cause and effect dynamics using dedicated encoders. By separately encoding the cause and effect components of the video and then combining them to generate the final caption, CEN is able to better understand and articulate the causal-temporal narrative. The primary contributions of our work are:
@@
  • Introducing for the first time CTN captions benchmark with high-quality, coherent causal-temporal narrative captions, automatically evaluated, to address the lack of coherent representations in video captioning benchmarks.
@@
Our work marks a significant step forward in video captioning research by explicitly addressing the challenges of causal-temporal narrative understanding. The CTN captions benchmark and the CEN architecture provide a comprehensive framework for learning and generating contextually relevant and temporally coherent video descriptions. This framework has the potential to drive the development of more sophisticated video understanding systems and enhance various applications, such as video summarization, content retrieval, and human-robot interaction.
@@ 2.1 Benchmarks
MSVD [7] is a benchmark focused on human activities, providing a platform for evaluating video captioning models. The captions in MSVD often describe the observable actions without delving into the underlying motivations or the cause-effect relationships between the events. MSR-VTT [55] is a large-scale benchmark with diverse video content, encompassing a wide range of topics and genres. The captions often focus on describing the observable content without capturing the causal links between the events or the temporal progression of the narrative. As a result, models trained on MSVD and MSR-VTT may struggle to generate descriptions that accurately reflect the causal and temporal dynamics in the videos.
While recent benchmarks like NExT-QA [54] and EgoSchema [30] have made strides in incorporating causal and temporal reasoning, they focus on specific aspects of video understanding, such as question-answering and long-form video comprehension. NExT-QA [54] introduces multi-choice and open-ended question-answering tasks that target causal and temporal action reasoning but do not address the generation of holistic video descriptions. EgoSchema [30], on the other hand, emphasizes long-form video understanding and temporal reasoning but does not explicitly focus on causal-temporal narrative for video captioning.
Similarly, efforts like VCR [60], V2C [12], and Motivation [49] integrate causality into their analysis of visual description or question-answering, relying heavily on commonsense reasoning for generating predictions. VCR [60] focuses on visual commonsense reasoning, V2C [12] aims to generate commonsense descriptions for video captioning, and Motivation [49] explores the prediction of motivations behind actions in videos. However, these works primarily rely on commonsense reasoning and do not delve into the causal and temporal structures underpinning video narratives. We introduce a new benchmark CTN specifically designed to capture and evaluate causal-temporal narrative in video captioning.
@@ 2.2 Video Captioning
Video captioning techniques have evolved from LSTM-based [13, 42, 32] frameworks to the latest designs using SOTA GNNs [17, 62, 34] and Transformers [51, 24, 58], with a focus on enhancing the complexity of captions through the injection of multimodal data. Despite these advancements, current architectures often struggle to capture the intricate temporal sequences and causal relationships in video storytelling. To bridge this gap, video captioning can benefit from cross-fertilization with ideas and strategies developed in related fields, such as action recognition [43, 50, 21, 53, 8, 14, 35, 38, 2, 36, 61], event localization [45, 25, 11, 26], and question-answering [1, 18, 39, 15, 59, 22, 40, 33]. The integration of causal reasoning [27, 57] has shown promise in enhancing the ability of neural networks to discern causal relationships, leading to improved performance in image captioning [27] and visual question answering [57]. However, current SOTA models still struggle to effectively handle the narrative complexity in videos, particularly in terms of causal and temporal progression. To address this limitation, there is a need to develop new video captioning models that can overcome the shortcomings of current frameworks.
@@ 3 Method
3.1 CTN Captions Benchmark
To address the challenges of existing benchmarks in representing causal and temporal relationships within video content, we propose an approach that harnesses the potential of LLMs through the few-shot prompting technique [5]. Our method generates CTN captions without the need for model fine-tuning, leveraging the inherent generative capabilities of LLMs to produce captions that encapsulate causal-temporal narrative structures. Figure 2 shows the CTN caption generation pipeline, which consists of two key steps: prompt design, LLM-based caption generation and evaluation.
Prompt Design: The prompt design step is crucial in guiding the LLM to generate causal-temporal narrative captions. The input to the CTN caption generation are the original video captions from existing benchmarks. We design a prompt that include a small set of carefully selected example captions, illustrating the desired output structure and highlighting the key aspects of causal-temporal narrative. We design Prompt 1 which guides LLM to generate accurate and concise causal-temporal narrative. The prompt sets clear requirements for grounding the captions in the provided descriptive context, maintains temporal consistency, and avoids unsupported details or conclusions. The illustrative examples demonstrate the desired format, facilitating the generation of plainly written, length-constrained Cause and Effect statements that effectively capture the video’s causal-temporal narrative. By explicitly outlining instructions and constraints, the carefully designed prompt steers the LLM’s generative capabilities towards producing contextually relevant and narratively coherent CTN captions. Further details are provided in Appendix B and C.
CTN caption Generation and Evaluation: We send Prompt 1 into the Mixtral of Experts LLM [20], which generates CTN captions that accurately reflect the causal-temporal narrative in the video. The LLM’s advanced natural language understanding and generation capabilities, combined with the few-shot prompting approach, enable the production of captions that encapsulate complex causal and temporal relationships. To ensure the quality and relevance of the generated CTN captions, we employ the EMScore [41] metric for evaluation. EMScore directly measures the consistency between a caption and the video content, and has been shown to be more effective than other metrics [10] in evaluating video relevance without referenced captions. We set a threshold ($\theta$ = 0.2) for the EMScore value, indicating adequate relevance to the video. Captions with an EMScore above the $\theta$ are kept, while those below the $\theta$ are discarded, and the LLM generates a new caption. This iterative refinement process continues until a caption meets the EMScore threshold, ensuring that the final captions efficiently describe the relevant events in the video. Further details are provided in Appendix D.
@@ 3.2 Proposed Cause-Effect Network (CEN)
Existing SOTA video captioning methods based on LSTMs[32], GNNs[17], and Transformers[51] lack dedicated mechanisms to explicitly model the intricate causal-temporal narrative sequences. Our proposed CEN employs a two-stage approach to address this limitation. In Stage 1 (Section 3.2.1), CEN learns specialized cause and effect video representations using separate encoders trained on the corresponding portions of the CTN captions. The input is raw video frames, and the output is the learned cause and effect features extracted from the encoders. The learned video features from Stage 1 are then passed to Stage 2 (Section 3.2.2), which utilizes a sequence-to-sequence transformer-based network with two new encoders to generate the final captions. The parameters of stage 1 encoders are frozen at this stage. The cause and effect video features are separately encoded and then concatenated before being input to the decoder. This allows the model to synthesize the specialized representations learned in Stage 1 and generate captions that accurately capture the causal and temporal relationships in the videos.

### v2

@@ Abstract
Existing video captioning benchmarks and models lack coherent representations of causal-temporal narrative, which is sequences of events linked through cause and effect, unfolding over time and driven by characters or agents. This lack of narrative restricts models’ ability to generate text descriptions that capture the causal and temporal dynamics inherent in video content. To address this gap, we propose NarrativeBridge, an approach comprising of: (1) a novel Causal-Temporal Narrative (CTN) captions benchmark generated using a large language model and few-shot prompting, explicitly encoding cause-effect temporal relationships in video descriptions, evaluated automatically to ensure caption quality and relevance and validated through human evaluation; and (2) a dedicated Cause-Effect Network (CEN) architecture with separate encoders for capturing cause and effect dynamics independently, enabling effective learning and generation of captions with causal-temporal narrative. Extensive experiments demonstrate that CEN significantly outperforms state-of-the-art models, including fine-tuned vision-language models, and is more accurate in articulating the causal and temporal aspects of video content than the second best model (GIT): 17.88 and 17.44 CIDEr on the MSVD and MSR-VTT datasets, respectively. Cross-dataset evaluations further showcase CEN’s strong generalization capabilities. The proposed framework understands and generates nuanced text descriptions with intricate causal-temporal narrative structures present in videos, addressing a critical limitation in video captioning. For project details, visit https://narrativebridge.github.io/.
@@ 1 Introduction
Video captioning aims to generate textual descriptions that capture the visual information and temporal dynamics in videos Venugopalan et al. (2015). Research has primarily focused on developing new methods to improve the accuracy of video captioning models on well-established benchmark datasets, MSR-VTT Xu et al. (2016) and MSVD Chen & Dolan (2011). State-of-the-art (SOTA) approaches Iashin & Rahtu (2020); Tian et al. (2019); Xu et al. (2017); Nadeem et al. (2023); Wang et al. (2022); Chen et al. (2023) proposed new architectures to better align the generated captions with the provided ground-truth. However, these efforts enhance models’ accuracy on existing evaluation criteria, without modifying the benchmark datasets or their ground truth annotations to address the lack of coherent representations of causal-temporal narrative - construction and interpretation of a sequence of events linked through cause and effect, unfolding over time and space, and often driven by entities (characters or agents) acting with intention Wilkens et al. (2003).
As discussed by Wilkens et al. (2003), causal relationships Granger (1969) are scenarios where one event (cause) directly influences the occurrence of another event (effect), while temporal understanding Berrevoets et al. (2023) involves recognizing the chronological order of events, a crucial component in establishing causality. Existing ground-truth captions in popular benchmarks lack causal, temporal and narrative information. Figure 1 illustrates the importance of causal-temporal narrative on the MSR-VTT dataset Xu et al. (2016). In the input video sequence, a car is first shown driving recklessly through an open field and flipping over, which represents the cause event. This is then followed by the effect event showing the severely damaged car, and subsequently, a group of guys starting to play beer pong, which represents the temporal sequence. The causal-temporal narrative would connect these events in a temporal order, linking the reckless driving and resulting car crash to the group’s subsequent behavior of playing beer pong. This example aligns with the definition of narrative provided by Wilkens et al. Wilkens et al. (2003), which emphasizes the role of cause-effect relationships, the perception of narrativity in videos, and highlights the temporal sequence of events that is crucial for understanding the causal-temporal narrative. Figure 1 highlights the limitations of the original captions in existing benchmark datasets, such as MSR-VTT Xu et al. (2016). The original captions focus on isolated events or actions, such as "a car flipping over" or "a monster truck flips on its side," lacking contextual narrative and causal-temporal relationships between events. Consequently, models trained on these captions suffer from the same limitations.
@@
To bridge this gap, we introduce NarrativeBridge, a novel framework encompassing a new benchmark and architecture tailored for causal-temporal narrative learning in video captioning. Our Causal-Temporal Narrative (CTN) captions benchmark leverages a large language model (LLM) and few-shot prompting to generate enhanced video descriptions that explicitly encode cause-effect relationships and temporal sequences, as exemplified in Figure 1. Our CTN captions benchmark, generated using the approach detailed in Section 3, establishes a clear connection between the cause (reckless driving) and the effect (damaged car and subsequent behavior of the group). By generating a single, coherent caption that encapsulates the causal-temporal narrative of the video, our CTN captions benchmark enables models to better understand and articulate the causality, sequence, and significance of events within the broader video context Wilkens et al. (2003). This approach addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions. To ensure the quality and relevance of the generated captions, we employ an automatic evaluation framework that compares the CTN captions with the video content, keeping or discarding them based on a score threshold. Additionally, we conduct a human evaluation study that further validates the high quality of our CTN captions, demonstrating their accuracy, temporal coherence, and relevance. This CTN captions benchmark addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions.
@@
Existing SOTA video captioning methods that use different architectures such as LSTM Nadeem et al. (2023), GNN Hendria et al. (2023), and Transformer Wang et al. (2022), struggle to effectively learn causal-temporal narrative from the CTN captions. These architectures are designed to capture the overall semantics in videos but lack dedicated mechanisms to explicitly model the cause-effect relationships and temporal sequences. As a result, the captions generated by these methods fail to articulate the complex causal-temporal dynamics in the videos, as demonstrated in results (Section 4.4). To address this challenge, we propose the Cause-Effect Network (CEN), an architecture specifically designed to capture cause and effect dynamics using dedicated encoders. By separately encoding the entire video through the cause and effect encoders, without requiring explicit cause-effect segmentation, and then combining them to generate the final caption, CEN is able to better understand and articulate the causal-temporal narrative. The primary contributions of our work are:
@@
  • Introducing for the first time CTN captions benchmark with high-quality, coherent causal-temporal narrative captions, automatically evaluated and humanly validated, to address the lack of coherent representations in video captioning benchmarks.
@@
In addition, our work demonstrates strong generalization capabilities through cross-dataset evaluations and outperforms even fine-tuned SOTA vision-language models such as VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) in generating causal-temporal narratives. Our work marks a significant step forward in video captioning research by explicitly addressing the challenges of causal-temporal narrative understanding. The CTN captions benchmark and the CEN architecture provide a comprehensive framework for learning and generating contextually relevant and temporally coherent video descriptions.
@@ 2.1 Benchmarks
MSVD Chen & Dolan (2011) is a benchmark focused on human activities, providing a platform for evaluating video captioning models. The captions in MSVD often describe the observable actions without delving into the underlying motivations or the cause-effect relationships between the events. MSR-VTT Xu et al. (2016) is a large-scale benchmark with diverse video content, encompassing a wide range of topics and genres. The captions often focus on describing the observable content without capturing the causal links between the events or the temporal progression of the narrative. As a result, models trained on MSVD and MSR-VTT may struggle to generate descriptions that accurately reflect the causal and temporal dynamics in the videos.
While recent benchmarks like NExT-QA Xiao et al. (2021) and EgoSchema Mangalam et al. (2024) have made significant strides in incorporating causal and temporal reasoning in video understanding, they focus primarily on question-answering tasks rather than generating comprehensive causal-temporal narratives. NExT-QA introduces multi-choice and open-ended question-answering tasks that target causal action reasoning but do not address the generation of holistic video descriptions. It focuses on specific question-answer pairs that often target single events or actions. In contrast, our CTN captions provide a more comprehensive narrative that captures the causal and temporal relationships across the entire video sequence. It means that our CTN benchmark differs significantly by focusing on generating comprehensive causal-temporal narratives, capturing broader temporal relationships within a single caption, and providing a more holistic view of video content (see Appendix A.7 for a detailed comparison). EgoSchema Mangalam et al. (2024), on the other hand, emphasizes long-form video understanding and temporal reasoning but does not explicitly focus on causal-temporal narrative for video captioning.
Similarly, efforts like VCR Zellers et al. (2019), V2C Fang et al. (2020), and Motivation Vondrick et al. (2016) integrate causality into their analysis of visual description or question-answering, relying heavily on commonsense reasoning for generating predictions. VCR Zellers et al. (2019) focuses on visual commonsense reasoning, V2C Fang et al. (2020) aims to generate commonsense descriptions for video captioning, and Motivation Vondrick et al. (2016) explores the prediction of motivations behind actions in videos. However, these works primarily rely on commonsense reasoning and do not delve into the causal and temporal structures underpinning video narratives. We introduce a new benchmark CTN specifically designed to capture and evaluate causal-temporal narrative in video captioning. Our approach goes beyond existing benchmarks by explicitly modeling causal-temporal narrative in a single, coherent caption, enabling a more comprehensive understanding of video content.
@@ 2.2 Video Captioning
Video captioning techniques have evolved from LSTM-based Gao et al. (2017); Song et al. (2017); Nadeem et al. (2023) frameworks to the latest designs using SOTA GNNs Hendria et al. (2023); Zhang et al. (2020); Pan et al. (2020) and Transformers Wang et al. (2022); Lin et al. (2022); Yang et al. (2023), with a focus on enhancing the complexity of captions through the injection of multimodal data. Despite these advancements, current architectures often struggle to capture the intricate temporal sequences and causal relationships in video storytelling. To bridge this gap, video captioning can benefit from cross-fertilization with ideas and strategies developed in related fields, such as action recognition Sun et al. (2022); Wang et al. (2016); Kazakos et al. (2019); Xiao et al. (2020); Chen & Ho (2022); Gao et al. (2020); Panda et al. (2021); Sardari et al. (2023); Alfasly et al. (2022); Planamente et al. (2021); Zhang et al. (2022), event localization Tian et al. (2018); Lin et al. (2019); Duan et al. (2021); Lin et al. (2021), and question-answering Alamri et al. (2019); Hori et al. (2019); Schwartz et al. (2019); Geng et al. (2021); Yun et al. (2021); Li et al. (2022a); Shah et al. (2022); Nadeem et al. (2024). The integration of causal reasoning Liu et al. (2022); Xue et al. (2023) has shown promise in enhancing the ability of neural networks to discern causal relationships, leading to improved performance in image captioning Liu et al. (2022) and visual question answering Xue et al. (2023). However, current SOTA models still struggle to effectively handle the narrative complexity in videos, particularly in terms of causal and temporal progression.
Recent advancements in vision-language models (VLMs) such as VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) have shown promising results in various video understanding tasks. However, as our experiments show (see Section 4.4), even these advanced models struggle with generating accurate causal-temporal narratives. To address this limitation, there is a need to develop new video captioning models that can overcome the shortcomings of current frameworks. This also underscores the need for specialized architectures like our proposed CEN that are explicitly designed to capture and generate causal-temporal narratives in videos.
@@ 3 Method
3.1 CTN Captions Benchmark
To address the challenges of existing benchmarks in representing causal and temporal relationships within video content, we propose an approach that harnesses the potential of LLMs through the few-shot prompting technique Brown et al. (2020). Our method generates CTN captions without the need for model fine-tuning, leveraging the inherent generative capabilities of LLMs to produce captions that encapsulate causal-temporal narrative structures. Figure 2 shows the CTN caption generation pipeline, which consists of two key steps: prompt design, LLM-based caption generation and evaluation.
Prompt Design: The prompt design step is crucial in guiding the LLM to generate causal-temporal narrative captions. The input to the CTN caption generation are the original video captions from existing benchmarks. We design a prompt that include a small set of carefully selected example captions, illustrating the desired output structure and highlighting the key aspects of causal-temporal narrative. We design Prompt 1 which guides LLM to generate accurate and concise causal-temporal narrative. The prompt sets clear requirements for grounding the captions in the provided descriptive context, maintains temporal consistency, and avoids unsupported details or conclusions. The illustrative examples demonstrate the desired format, facilitating the generation of plainly written, length-constrained Cause and Effect statements that effectively capture the video’s causal-temporal narrative. By explicitly outlining instructions and constraints, the carefully designed prompt steers the LLM’s generative capabilities towards producing contextually relevant and narratively coherent CTN captions. Further details are provided in Appendix A.2 and A.3.
@@ 3.1 CTN Captions Benchmark
CTN Caption Generation and Automatic Evaluation: We send Prompt 1 into the Mixtral of Experts LLM Jiang et al. (2024), which generates CTN captions that accurately reflect the causal-temporal narrative in the video. The LLM’s advanced natural language understanding and generation capabilities, combined with the few-shot prompting approach, enable the production of captions that encapsulate complex causal and temporal relationships. To ensure the quality and relevance of the generated CTN captions, we employ the EMScore Shi et al. (2022) metric for evaluation. EMScore directly measures the consistency between a caption and the video content, and has been shown to be more effective than other metrics de Souza Inácio & Lopes (2023) in evaluating video relevance without referenced captions. We set a threshold ($\theta$ = 0.2) for the EMScore value, indicating adequate relevance to the video. Captions with an EMScore above the $\theta$ are kept, while those below the $\theta$ are discarded, and the LLM generates a new caption. This iterative refinement process continues until a caption meets the EMScore threshold, ensuring that the final captions efficiently describe the relevant events in the video. Further details are provided in Appendix A.4.
Human Evaluation: To further validate the quality of our CTN captions, we conduct a human evaluation study. Five independent domain experts evaluate 100 randomly sampled captions (50 each from MSRVTT-CTN and MSVD-CTN) on three criteria: causal accuracy - the degree to which the caption correctly identifies and describes cause-effect relationships; temporal coherence - the extent to which the caption accurately represents the sequence of events; and relevance - how well the caption reflects the overall content. Each criterion is rated on a 5-point Likert scale. Following are the results: the overall mean quality score of 4.8 ($\sigma$ = 0.40) indicates high performance across all criteria; 93% of captions receive scores of 4 or higher across all three dimensions, with 84% achieving perfect scores. To ensure reliability across raters, we calculate the intraclass correlation coefficient (ICC). The ICC for absolute agreement among raters is 0.87 (95% CI: 0.83-0.91), indicating high inter-rater reliability. These results validate our automatic generation and evaluation process, demonstrating that our CTN captions benchmark provides high-quality, coherent representations of causal-temporal narratives in videos. Further details are provided in Appendix A.5.
@@ 3.2 Proposed Cause-Effect Network (CEN)
Existing SOTA video captioning methods based on LSTMsNadeem et al. (2023), GNNsHendria et al. (2023), and TransformersWang et al. (2022) lack dedicated mechanisms to explicitly model the intricate causal-temporal narrative sequences. Our proposed CEN employs a two-stage approach to address this limitation. In Stage 1 (Section 3.2.1), CEN learns specialized cause and effect video representations using separate encoders trained on the corresponding portions of the CTN captions. The input is raw video frames from the entire video without requiring explicit segmentation of cause-effect regions, and the output is the learned cause and effect features extracted from the encoders. The learned video features from Stage 1 are then passed to Stage 2 (Section 3.2.2), which utilizes a sequence-to-sequence transformer-based network with two new encoders to generate the final captions. The parameters of stage 1 encoders are frozen at this stage. The cause and effect video features are separately encoded and then concatenated before being input to the decoder. This allows the model to synthesize the specialized representations learned in Stage 1 and generate captions that accurately capture the causal and temporal relationships in the videos.